What Is The Equation Of The 2 Points (-4,-4)(3,-2)
Answer by rfer(12730)

slope
m=(-2--4)/(3--4)
m=2
----------------
y intercept
y=mx+b
-2=2*3+b
-2=6+b
-2-6=b
-8=b
-----------------
y=2x-8